If a boater can travel 22 miles downstream in the same time it takes to travel 16 miles upstream, with a current of 4 mi/h, what is the rate of the boat instill water? I know the formula for distance, but I can't seem to set up the equation correctly. Please help me!
Answer by jrfrunner(365)

distance = rate*time or speed *time
downstream the speed is Boat + current speed : B+C
upstream the speed is Boat -current speed: B-C
--
22miles = (B+C)*t=Bt+Ct
16miles = (B-C)*t=Bt-Ct
--
since C (current speed)=4
22miles =(B+4)*t=B*t+4t
16 miles =(B-4)*t=B*t-4*t
--
add these two equations
--
38 = 2Bt
solve for t: t=38/(2B)=19/B
subsitute t into one of the orginal equations
---
22 = B(19/B) + 4(19B)
22=19+76/B
can you handled it from here, by solving for B (the boat speed in still water)?